TiMoThYnICee:

Write 26/20 as a percent

TiMoThYnICee:

26/20 converted to percent would be 130%. 26/20 = 13/10 Divide 13 and 10 and you get 1.3 1.3 * 100 = 130%.

TiMoThYnICee:

that's the answer hope this helps

Vijeya3:

who did ya post this for lol?

TiMoThYnICee:

lol I knew this answer already I'm doing this for everybody pal

Vijeya3:

oh...not to increase ur SS?x'D

TiMoThYnICee:

xD no I want to help others idc about my ss

TiMoThYnICee:

these are questions I already answered on brainly, if people don't like brainly or want to visit it or cant, than I'm bringing my answers here for everyone

Arcadiouse:

But you close it right after...

TiMoThYnICee:

yeah so I can ask the next question of course

Vijeya3:

loll m proud of ya buddyx'D

Arcadiouse:

Yeah ok.